### v1

# A closer look at some cyclic semifields

Abstract.
We show that different choices of generators of the Galois group of $\mathbb{F}_{q^{n}}/\mathbb{F}_{q}$ ($n\geq 3$) produce non-isotopic cyclic semifields $\mathbb{F}_{q^{n}}[t;\sigma]/\mathbb{F}_{q^{n}}[t;\sigma](t^{m}-a)$ when $m=n$: for $n\geq m$, there are $\varphi(n)$ non-isomorphic classes of Sandler semifields $\mathbb{F}_{q^{n}}[t;\sigma]/\mathbb{F}_{q^{n}}[t;\sigma](t^{m}-a)$, one class for each generator $\sigma$ of ${\rm Gal}(\mathbb{F}_{q^{n}}/\mathbb{F}_{q})$ involved in their construction, where $\varphi$ is the Euler function. We prove that when $n=m$, two Sandler semifields constructed from different generators $\sigma_{1}$ and $\sigma_{2}$ of ${\rm Gal}(\mathbb{F}_{q^{n}}/\mathbb{F}_{q})$ are not isotopic. Hence when $n=m$ there are $\varphi(m)$ non-isotopic classes of these semifields, each class belonging to one choice of generator. We then present a full parametrization of the non-isomorphic Sandler semifields of order $q^{m^{2}}$ with center $\mathbb{F}_{q}$, and nuclei $\mathbb{F}_{q^{m}}$, when $m$ is prime and $\mathbb{F}_{q}$ contains a primitive $m$th root of unity. Since for $m=n$, two Sandler semifields constructed from the same generator are isotopic if and only if they are isomorphic, this parametrizes these Sandler semifields up to isotopy, and thus the corresponding non-Desarguesian projective planes as well. Most of our results are proved in all generality for any cyclic Galois field extension.
Key words and phrases:
Semifields, isotropy, isometry, cyclic semifields
2020 Mathematics Subject Classification:
17A35, 17A99
1. Introduction
Petit algebras were introduced in [Pe1966, Pet1968] and then rediscovered over finite fields in [LaShe2013]. They yield a large class of semifields, i.e. of unital nonassociative division algebras over a finite field. Over finite base fields, Petit division algebras are isotopic to cyclic semifields (also called Jha-Johnson semifields) introduced in [JJ989]. Over general base fields, Petit algebras, denoted by $S_{f}=D[t;\sigma,\delta]/D[t;\sigma,\delta]f$, are constructed using a skew polynomial ring $D[t;\sigma,\delta]$, where $D$ is a unital associative ring, and some monic polynomial $f\in D[t;\sigma,\delta]$.
Let $K/F$ be a cyclic field extension of degree $n$ with Galois group ${\rm Gal}(K/F)$, and let $\sigma$ and $\sigma^{\prime}$ be two different generators of ${\rm Gal}(K/F)$. The question whether two Petit algebras $K[t;\sigma]/K[t;\sigma]f$ and $K[t;\sigma^{\prime}]/K[t;\sigma^{\prime}]f^{\prime}$ that are constructed employing different generators $\sigma$ and $\sigma^{\prime}$ of ${\rm Gal}(K/F)$ yield isotopic algebras was raised for the first time in [LaShe2013, Remark 6] in the context of semifields. It is known that every semifield $K[t;\sigma]/K[t;\sigma]f$ is isotopic to a semifield $K[t;\sigma^{-1}]/K[t;\sigma^{-1}]\bar{f}$, where $\bar{f}$ is the reciprocal of $f$ [KaLieb2008].
In this paper, we show that in certain cases, different choices of generators $\sigma_{1}$ and $\sigma_{2}$ of ${\rm Gal}(K/F)$ can produce non-isotopic Petit algebras. This is a consequence of our First Main Theorem 3.3:
Theorem.
Let $n\geq 3$ and let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of the Galois group $\mathrm{Gal}(K/F)$. Suppose that $n\geq m$. Let $t^{m}-a_{1}\in K[t;\sigma_{1}]$ and $t^{m}-a_{2}\in K[t;\sigma_{2}]$, $a_{i}\in K\smallsetminus F$, both be not right-invariant. Then
$K[t;\sigma_{1}]/K[t;\sigma_{1}](t^{m}-a_{1})\not\cong K[t;\sigma_{2}]/K[t;\sigma_{2}](t^{m}-a_{2}).$
For $m=n$, the algebras $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ are also denoted by $(K/F,\sigma,a)$ and called nonassociative cyclic algebras of degree $m$, as they can be seen as canonical generalizations of associative cyclic algebras; for $a\in F^{\times}$, the algebra $(K/F,\sigma,a)$ is a classical cyclic algebra of degree $m$ as defined in [KMRT, p. 414].
When $F=\mathbb{F}_{q}$ is a finite field, $n\geq m$, and both $t^{m}-a_{i}\in\mathbb{F}_{q^{n}}[t;\sigma_{i}]$, $i\in\{1,2\}$, are irreducible, then $\mathbb{F}_{q^{n}}[t;\sigma_{1}]/\mathbb{F}_{q^{n}}[t;\sigma_{1}](t^{m}-a_{1})$ and $\mathbb{F}_{q^{n}}[t;\sigma_{2}]/\mathbb{F}_{q^{n}}[t;\sigma_{2}](t^{m}-a_{2})$ are non-isomorphic semifields (Corollary 4.3). Semifields of the type $\mathbb{F}_{q^{n}}[t;\sigma]/\mathbb{F}_{q^{n}}[t;\sigma](t^{m}-a)$, $n\geq m$, are the opposite algebras (i.e., the dual semifields) of Sandler’s semifields [San1962]. In particular, hence there are $\varphi(m)$ ($\varphi$ the Euler function) different classes of non-isomorphic Sandler semifields that are the dual semifields of
$\mathbb{F}_{q^{m}}[t;\sigma_{i}]/\mathbb{F}_{q^{m}}[t;\sigma_{i}](t^{m}-a)=(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma,a)$
over a given base field $\mathbb{F}_{q}$, one class for each choice of generator $\sigma_{i}$ of the Galois group of $\mathbb{F}_{q^{m}}/\mathbb{F}_{q}$.
When $n=m$, two Sandler semifields constructed with the same $\sigma$ are isotopic algebras if and only if they are isomorphic [San1962, Theorem 9]. We generalize this result in our other Second Main Theorem 4.1:
Theorem.
Let $K/F$ be a cyclic Galois field extension, $m\geq 3$, and $\sigma_{1}$, $\sigma_{2}$ two generators of ${\rm Gal}(K/F)$. Let $(K/F,\sigma_{1},a_{1})$ and $(K/F,\sigma_{2},a_{2})$ be two proper nonassociative algebras. If $(K/F,\sigma_{1},a_{1})$ and $(K/F,\sigma_{2},a_{2})$ are isotopic, then $\sigma_{1}=\sigma_{2}$.
Hence for fixed $m\geq 3$, two proper semifields $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{1},a_{1})$ and $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{2},a_{2})$ are isotopic if and only if they are isomorphic.
When $m$ is prime and $\mathbb{F}_{q}$ contains a primitive root of unity, we count the proper nonassociative cyclic algebras (i.e., those Sandler semifields where $m=n$) up to isomorphism and thus up to isotropy, and give an explicit parametrization of the isomorphism (and thus, respectively, isotopism) classes of these algebras corresponding to a fixed $F=\mathbb{F}_{q}$. Since different choices of generators $\sigma_{1}$ and $\sigma_{2}$ of ${\rm Gal}(\mathbb{F}_{q^{m}}/\mathbb{F}_{q})$ produce non-isomorphic nonassociative cyclic algebras of degree $m$, there are more non-isomorphic Sandler semifields than previously counted.
The paper is organized as follows. After the preliminaries in Section 2, Section 3 contains the work towards our First Main Theorem, cited above. In Section 4, we prove our Second Main Theorem 4.1, count the number of Sandler semifields that are opposite algebras of nonassociative cyclic algebras of prime degree $m$ (Theorem 4.6) and give a parametrization of the algebras $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{1},a_{1})$ (Theorems 5.1 and 5.2) under the assumption that the base field $F$ contains a primitive $m$th root of unity in Section 5. Since two proper semifields coordinatize the same non-Desarguesian projective plane if and only if they are isotopic, we have thus also obtained a parametrization of the non-Desarguesian projective planes related to $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{1},a_{1})$.
We finish with a brief outlook on some open questions and what our results mean for linear codes in Section 6.
2. Preliminaries
We use $R^{\times}$ to denote the group of invertible elements of a unital associative ring $R$, and denote by $(R^{\times})^{n}$ the subgroup $\{x^{n}\mid x\in R^{\times}\}$. Let $F$ be a field.
2.1. Nonassociative algebras
An $F$-vector space $A$ is an algebra over $F$ if there exists an $F$-bilinear map $A\times A\to A$, $(x,y)\mapsto x\cdot y$. We denote this multiplication in $A$ simply by the juxtaposition $xy$. An algebra $A$ is called unital if there is an element in $A$, denoted by 1, such that $1x=x1=x$ for all $x\in A$. We will only consider unital finite-dimensional nonzero algebras. If $A$ is a unital associative algebra, then we denote the group of invertible elements in $A$ by $A^{\times}$.
Define the associator of $A$ by $[x,y,z]=(xy)z-x(yz)$. The left, middle and right nucleus of $A$ are defined as ${\rm Nuc}_{l}(A)=\{x\in A\,|\,[x,A,A]=0\}$, ${\rm Nuc}_{m}(A)=\{x\in A\,|\,[A,x,A]=0\}$ and ${\rm Nuc}_{r}(A)=\{x\in A\,|\,[A,A,x]=0\}$, respectively. These are associative subalgebras of $A$ and their intersection ${\rm Nuc}(A)=\{x\in A\,|\,[x,A,A]=[A,x,A]=[A,A,x]=0\}$ is the nucleus of $A$. The center of $A$ is ${\rm C}(A)=\{x\in A\,|\,x\in\text{Nuc}(A)\text{ and }xy=yx\text{ for all }y\in A\}$.
An algebra $A\not=0$ is called a division algebra if for any $a\in A$, $a\not=0$, the left multiplication with $a$, $L_{a}(x)=ax$, and the right multiplication with $a$, $R_{a}(x)=xa$, are bijective. If it is not clear which algebra structure on $V$ is considered, we denote the left and right multiplication with $a\in V$ by $L_{A}(a)$ and $R_{A}(a)$. If $A$ has finite dimension over $F$, $A$ is a division algebra if and only if $A$ has no zero divisors [Sch1995, pp. 15, 16].
We call a nonassociative algebra a proper nonassociative algebra, if it is not associative. A semifield is a unital nonassociative division algebra over a finite field. A semifield that is not associative is called a proper semifield.
We denote the set of non-unital algebra structures on an $F$-vector space $V$ by ${\rm Alg}(V)$. Given $A\in{\rm Alg}(V)$, we write $xAy$ for the product of $x,y\in V$ in the algebra $A$ instead of simply juxtaposition, when the algebra multiplication is not ad hoc clear from the context (subsequently, usually it will be obvious). For $f,g,h\in{\rm Gl}(V)$ the algebra $A^{(f,g,h)}$ is called an isotope of $A$, and is defined as the vector space $V$ together with the new multiplication
$xA^{(f,g,h)}y=h(f(x)Ag(y))$
for $x,y\in V$. Two algebras $A,A^{\prime}\in{\rm Alg}(V)$ are called isotopic, if $xAy=h(f(x)A^{\prime}g(y))$ for all $x,y\in V.$ If $f=g=h^{-1}$ then $A\cong A^{\prime}$. We know that $(A^{(f,g)})^{op}=(A^{op})^{(f,g)}$. If $A$ is a division algebra then $A^{(f,g,h)}$ is a division algebra.
2.2. Twisted polynomial rings
Let $K/F$ be a cyclic Galois field extension of degree $n$ with Galois group $G$ generated by $\sigma$.
The twisted polynomial ring $R=K[t;\sigma]$ is the ring of polynomials $\{a_{0}+a_{1}t+\dots+a_{m}t^{m}\,|\,a_{i}\in K\}$, with term-wise addition and multiplication given by the rule $ta=\sigma(a)t$ for all $a\in K$ [Ore1933]; see also [Jac1996, Ch I]. The constant nonzero polynomials $K^{\times}$ are the units of $R$.
For $f=a_{0}+a_{1}t+\dots+a_{m}t^{m}\in R$ with $a_{m}\not=0$ define ${\rm deg}(f)=m$ and put ${\rm deg}(0)=-\infty$. Then ${\rm deg}(fg)={\rm deg}(f)+{\rm deg}(g).$ An element $f\in R$ is called irreducible if it is not a unit and it has no proper factors, i.e. there do not exist $g,h\in R$, neither a unit, such that $f=gh$. There exists a right division algorithm in $R=K[t;\sigma]$: for all $g,f\in R$, $f\neq 0$, there exist unique $r,q\in R$ such that ${\rm deg}(r)<{\rm deg}(f)$ and $g=qf+r$ [Jac1996, §1.1]. (Our terminology is the one used by Petit [Pe1966] and different from Jacobson’s [Jac1996], who calls what we call right a left division algorithm and vice versa.)
2.3. Petit algebras, nonassociative cyclic algebras and Sandler’s semifields
Let $f\in R=K[t;\sigma]$ have degree $m>1$ and let ${\rm mod}_{r}f$ denote the remainder of right division by $f$. The vector space $R_{m}=\{g\in K[t;\sigma]\,|\,{\rm deg}(g)<m\}$ together with the multiplication $g\circ h=gh\,\,{\rm mod}_{r}f$, where the right hand side is the remainder obtained after dividing $gh$ on the right by $f$, is a unital nonassociative algebra over $F$ of dimension $mn$ denoted by $R/Rf$ or $S_{f}$, and called a Petit algebra.
When ${\rm deg}(g)+{\rm deg}(h)<m$, the multiplication of $g$ and $h$ in $R/Rf$ is the same as the multiplication of $g$ and $h$ in $R$. The algebra $R/Rf$ is associative if and only if $f$ is right-invariant. In that case, $R/Rf$ is the usual quotient algebra $K[t;\delta]/(f)$ [Pe1966, (7),(9), (10)].
If $A=K[t;\sigma]/K[t;\sigma]f$ is a proper Petit algebra, then $C(A)=F$,
${\rm Nuc}_{l}(A)={\rm Nuc}_{m}(A)=K$
and
${\rm Nuc}_{r}(A)=\{g\in S_{f}\,|\,fg\in Rf\}$
is the eigenspace of $f$ [Pe1966].
In this paper we will investigate the case when $f(t)=t^{m}-a\in K[t;\sigma]$ is not right-invariant, which is the case if and only if $a\in K\smallsetminus F$ [Pe1966]. The algebra $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ is a division algebra over $F$ if and only if $f(t)=t^{m}-a\in K[t;\sigma]$ is irreducible [Pe1966, (7)]. We also know that $A=K[t;\sigma]/K[t;\sigma](t^{m}-a)$ is a division algebra if and only if ${\rm Nuc}_{r}(A)$ is a division algebra, e.g. see [GoTor2014, Proposition 4]. Over finite base fields and for $m\leq n$, the opposite algebras of $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ are called Sandler’s semifields.
Theorem 2.1.
Let $A=K[t;\sigma]/K[t;\sigma](t^{m}-a)$ be a proper Petit algebra, i.e. $a\in K\setminus F$. Let $H=\{\tau\in G\,|\,\tau(a)=a\}$. Then $H=\langle\sigma^{s}\rangle$ for some integer $s$ such that $n=sr$. Put $E={\rm Fix}(\sigma^{s})$. Then
${\rm Nuc}_{r}(A)=K[t;\sigma^{s}]/K[t;\sigma^{s}](t^{m}-a)$
is an associative quotient algebra with center $E$. In particular, $K\subset{\rm Nuc}_{r}(A)$ and hence $K={\rm Nuc}(A)$. Moreover, if $n$ is prime then ${\rm Nuc}_{r}(A)=K$.
Proof.
A straightforward calculation shows that
${\rm Nuc}_{r}(A)=K\oplus Kt^{s}\oplus\dots\oplus Kt^{(r-1)s}.$
[…]
Let us turn to the question of when two such algebras are isomorphic.
Lemma 2.2.
Suppose $K,K^{\prime}$ are two distinct cyclic Galois extensions of a field $F$. Then for any generators $\sigma,\sigma^{\prime}$ of their respective Galois groups, and any $a\in K\smallsetminus F$, $a^{\prime}\in K^{\prime}\smallsetminus F$, we have
$K[t;\sigma]/K[t;\sigma](t^{m}-a)\not\cong K^{\prime}[t;\sigma^{\prime}]/K^{\prime}[t;\sigma^{\prime}](t^{m}-a^{\prime}).$
The proof is immediate, since isomorphic algebras will have the same left nuclei. In the rest of the paper, we will employ the following strong classification theorem which is a direct consequence of [Br2018, Theorem 4.9].
Theorem 2.3.
Let $K/F$ be a cyclic Galois extension of $F$ of degree $n$ and let $\sigma$ be a generator of $\mathrm{Gal}(K/F)$. Let $a,b\in K\smallsetminus F$. For $n\geq m$, we have
$K[t;\sigma]/K[t;\sigma](t^{m}-a)\cong K[t;\sigma]/K[t;\sigma](t^{m}-b)$
if and only if there exists some $\tau\in\mathrm{Gal}(K/F)$ and some $k\in K^{\times}$ such that
(2.1) $\tau(a)=\prod_{i=0}^{m-1}\sigma^{i}(k)b.$
For $n<m$, in Equation 2.1 holds for some $\tau\in\mathrm{Gal}(K/F)$ and some $k\in K^{\times}$, then
$K[t;\sigma]/K[t;\sigma](t^{m}-a)\cong K[t;\sigma]/K[t;\sigma](t^{m}-b).$
By applying the norm of the field extension to both sides of Equation 2.1, we obtain:
Corollary 2.4.
Let $K/F$ be a cyclic Galois extension of $F$ of degree $n\geq m$ and let $\sigma$ be a generator of $\mathrm{Gal}(K/F)$. If $a,b\in K\smallsetminus F$ such that
$N_{K/F}(a)\not\in(N_{K/F}(K^{\times}))^{m}N_{K/F}(b)$
then
$K[t;\sigma]/K[t;\sigma](t^{m}-a)\not\cong K^{\prime}[t;\sigma^{\prime}]/K[t;\sigma](t^{m}-b).$
Proof.
Suppose there exists some $\tau\in\mathrm{Gal}(K/F)$ and some $k\in K^{\times}$ such that $\tau(a)=\prod_{i=0}^{m-1}\sigma^{i}(k)b$. Apply $N_{K/F}$ on both sides to obtain that $N_{K/F}(\tau(a))=N_{K/F}(a)=(\prod_{i=0}^{m-1}N_{K/F}(\sigma^{i}(k))N_{K/F}(b)$ which implies the assertion observing that $N_{K/F}(\sigma^{i}(k))=N_{K/F}(k)$. ∎
Theorem 2.5.
[Br2018, Theorem 3.11] Suppose either that $m=2$ or $3$, or that $m\geq 5$ is a prime such that $F$ contains a primitive $m$th root of unity. Let $a\in K\smallsetminus F$. Then $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ is a division algebra if and only if
$\sigma^{m-1}(z)\cdots\sigma(z)z\not=a$
for all $z\in K$.
Corollary 2.6.
Suppose either that $m=2$ or $3$, or that $m\geq 5$ is a prime such that $F$ contains a primitive $m$th root of unity.
(i) Let $a\in K\smallsetminus F$ such that $N_{K/F}(a)\not\in N_{K/F}(K^{\times})^{m}$, then $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ is a division algebra.
(ii) If $F$ is a finite field and $N_{K/F}(a)\not\in(F^{\times})^{m}$, then $K[t;\sigma]/K[t;\sigma]f$ is a division algebra.
Proof.
Suppose there exists some $z\in K$, such that $\sigma^{m-1}(z)\cdots\sigma(z)z=a$. Apply $N_{K/F}$ to both sides if this equation, then
$N_{K/F}(\sigma^{m-1}(z)\cdots\sigma(z)z)=N_{K/F}(z)^{m}=N_{K/F}(a).$
This yields the assertion. ∎
We note that $n$ need not be equal to $m$ here.
For $m=n$, the unital central nonassociative algebra $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ is called a nonassociative cyclic algebra of degree $n$ and denoted by $(K/F,\sigma,a)$ (cf. [Ste2014] who studies the opposite algebras, however). When $a\in F^{\times}$, the associative algebra $(K/F,\sigma,a)$ is a classical associative central simple cyclic algebra of degree $m$ [Jac1996, p. 19]. For $n=m$ we also have a stronger result than Corollary 2.6:
Theorem 2.7.
[Pe1966] Suppose either that $m=2$ or $3$, or that $m\geq 5$ is a prime such that $F$ contains a primitive $m$th root of unity. Let $a\in K\smallsetminus F$. Then $(K/F,\sigma,a)$ is a division algebra.
By Theorem 4.2, for $H=\{\tau\in G\,|\,\tau(a)=a\}$ there is some integer $s$ such that $n=sr$ and $H=\langle\sigma^{s}\rangle$. Then
${\rm Nuc}_{r}((K/F,\sigma,a))=(K/E,\sigma^{s},a)$
[…]
Corollary 2.8.
Let $F$ be a finite field. If $s>1$ then $(K/F,\sigma,a)$ is not a semifield.
Proof.
Over finite fields there are no cyclic algebras which are division algebras other than the field itself. So here the right nucleus is a cyclic algebra of degree $r>1$, hence has zero divisors and so does the algebra $(K/F,\sigma,a)$. ∎
3. Different generators of $\mathrm{Gal}(K/F)$ yield nonisomorphic Sandler algebras
Let $F$ be an arbitrary field such that $F$ admits a cyclic Galois extension $K$ of degree $n>2$ and let $\sigma$ be a generator of the Galois group of $K/F$. We will now generalize Theorem [NevPu2024, Theorem 5.1] which covers the case $n=m$ that occurs when studying nonassociative cyclic algebras $(K/F,\sigma,a)$ to the case that $n\geq m$, starting with some general observations.
Let $f,g\in R=K[t;\sigma]$ be not right-invariant. For two proper Petit algebras $R/Rf$ and $R/Rg$, every isomorphism $G:R/Rf\rightarrow R/Rg$ is induced by an isomorphism $G:K[t;\sigma]\rightarrow K[t;\sigma]$ which restricts to an automorphism $\tau\in{\rm Aut}(K)$ [Br2018, Theorem 4.9].
It is also straightforward to see that every isomorphism $G:K[t;\sigma]\rightarrow K[t;\sigma]$ canonically induces an isomorphism between the Petit algebras $K[t;\sigma]/K[t;\sigma]f$ and $K[t;\sigma]/K[t;\sigma]G(f)$, since $G|_{R_{m}}:R_{m}\rightarrow R_{m}$ and $G(g\circ_{f}h)=G(gh-sf)=G(h)G(g)-G(s)G(f)=G(g)\circ_{G(f)}G(h)$ for some $s\in R_{m}$.
Let now $K$ and $K^{\prime}$ be two cyclic Galois field extensions of $F$ of degree $n$ and $\mathrm{Gal}(K/F)=\langle\sigma_{1}\rangle$, $\mathrm{Gal}(K^{\prime}/F)=\langle\sigma_{2}\rangle$. Consider the twisted polynomial rings $R=K[t;\sigma_{1}]$ and $R^{\prime}=K^{\prime}[t;\sigma_{2}]$.
Proposition 3.1.
(i) Let $\tau:K\rightarrow K^{\prime}$ be an isomorphism. Then the map $G:K[t;\sigma_{1}]\rightarrow K^{\prime}[t;\sigma_{2}]$ defined via
$G(t)=\sum_{i=0}^{m-1}k_{i}t^{i},$
$k_{i}\in K$, extends $\tau$ to an isomorphism $G:K[t;\sigma_{1}]\rightarrow K^{\prime}[t;\sigma_{2}]$ if and only if $G(t)=kt$ for some $k\in K^{\times}$ and $\sigma_{2}\circ\tau=\tau\circ\sigma_{1}$.
(ii) Let $\sigma_{1}$ and $\sigma_{2}$ be two generators of ${\rm Gal}(K/F)$, and let $\tau\in{\rm Gal}(K/F)$. Then the map $G:K[t;\sigma_{1}]\rightarrow K[t;\sigma_{2}]$ defined via
$G(t)=\sum_{i=0}^{m-1}k_{i}t^{i},$
$k_{i}\in K$, extends $\tau$ to an isomorphism $G:K[t;\sigma_{1}]\rightarrow K[t;\sigma_{2}]$ if and only if $G(t)=kt$ for some $k\in K^{\times}$ and $\sigma_{1}=\sigma_{2}$.
Proof.
(i) By [Rim1978, Theorem 3], the map $G$ defined via $G(t)=\sum_{i=0}^{m-1}k_{i}t^{i}$, $k_{i}\in K$, extends $\tau$ to an isomorphism $G:K[t;\sigma_{1}]\rightarrow K^{\prime}[t;\sigma_{2}]$ if and only if
  1. (1)
    $\sigma_{2}(\tau(a))k_{i}=k_{i}\tau(\sigma_{1}^{i}(a))$ for all $a\in K$ and $i\in\{0,\dots m\},$
  2. (2)
    $k_{1}\in K^{\times},$
  3. (3)
    $k_{i}\text{ is nilpotent for all }i\in\{2,\dots m\}.$
Since $K$ and $K^{\prime}$ are fields they only have trivial nilpotent elements. This means (3) is the same as $k_{i}=0$ for all $i\in\{2,\dots m\}$. So in our situation the map $G$ defined via $G(t)=\sum_{i=0}^{m-1}k_{i}t^{i}$, $k_{i}\in K$, extends $\tau$ to an isomorphism $G:K[t;\sigma_{1}]\rightarrow K^{\prime}[t;\sigma_{2}]$ if and only if
  1. (1)
    $\sigma_{2}(\tau(a))k_{1}=k_{1}\tau(\sigma_{1}(a))$, $\sigma_{2}(\tau(a))k_{0}=k_{0}\tau(a)$ for all $a\in K$,
  2. (2)
    $k_{1}\in K^{\times},$
  3. (3)
    $k_{i}=0$ for all $i\in\{2,\dots m\}.$
Of course, $\sigma_{2}(\tau(a))k_{1}=k_{1}\tau(\sigma_{1}(a))$ here is equivalent to $\sigma_{2}\tau=\tau\sigma_{1}$ and hence $\sigma_{2}(\tau(a))k_{0}=k_{0}\tau(a)$ is the same as $k_{0}(\tau(\sigma_{1}(a))-a)=0$ for all $a\in K$, which means $k_{0}=0$, since $\sigma_{1}\not=id$. Thus (1), (2) and (3) are equivalent to $G(t)=kt$ for some $k\in K^{\times}$ and satisfies $\sigma_{2}\circ\tau=\tau\circ\sigma_{1}$.
(ii) follows from (i): $\tau$ extends to an isomorphism $G:K[t;\sigma_{1}]\rightarrow K[t;\sigma_{2}]$ if and only if $G(t)=kt$ for some $k\in K^{\times}$ and $\sigma_{2}(\tau(a))k=k\tau(\sigma_{1}(a))$ for all $a\in K$. This last condition means that we need
$\sigma_{2}\circ\tau=\tau\circ\sigma_{1}.$
If, as we do, we assume that $K^{\prime}=K$ is a cyclic Galois field extension, this is satisfied if and only if $\sigma_{1}=\sigma_{2}$. ∎
Lemma 3.2.
Let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of $\mathrm{Gal}(K/F)$. There exists $1<j<n$, with $(j,n)=1$, such that $\sigma_{1}=\sigma_{2}^{j}$. Let $f\in K[t;\sigma_{1}]$ and $g\in K[t;\sigma_{2}]$ be both not right-invariant of degree $m$, and suppose that $n\geq m$. Let
$H:K[t;\sigma_{1}]/K[t;\sigma_{1}]f\rightarrow K[t;\sigma_{2}]/K[t;\sigma_{2}]g$
be an $F$-algebra isomorphism between two proper nonassociative Petit algebras. Then $H$ restricts on $K$ to some $\tau\in{\rm Gal}(K/F)$ and
$H(t)=kt^{j}$
for some $k\in K^{\times}$.
Proof.
Suppose that $H:K[t;\sigma_{1}]/K[t;\sigma_{1}]f\rightarrow K[t;\sigma_{1}]/K[t;\sigma_{1}]g$; this is an isomorphism of $F$-vector spaces satisfying $H(uv)=H(u)H(v)$ for all $u,v\in K[t;\sigma_{1}]/K[t;\sigma_{1}]f$. Since $K[t;\sigma_{1}]/K[t;\sigma_{1}]f$ and $K[t;\sigma_{1}]/K[t;\sigma_{2}]g$ are not associative, they both have left nucleus $K$. Every isomorphism preserves the left nucleus, so $H$ restricts to some $\tau\in{\rm Gal}(K/F)$. Since the Galois group is cyclic, and thus abelian, $\tau$ commutes with $\sigma_{i}$, for $i\in\{1,2\}$. Suppose $H(t)=\sum_{i=0}^{m-1}k_{i}t^{i}$ for some $k_{i}\in K$. Then
(3.1) $H(tz)=H(t)H(z)=(\sum_{i=0}^{m-1}k_{i}t^{i})\tau(z)=\sum_{i=0}^{m-1}k_{i}\sigma_{2}^{i}(\tau(z))t^{i}$
and
(3.2) $H(tz)=H(\sigma_{1}(z)t)=\sum_{i=0}^{m-1}\tau(\sigma_{1}(z))k_{i}t^{i}$
for all $z\in K$. Comparing the coefficients of $t^{i}$ in (3.1) and (3.2) we obtain
(3.3) $k_{i}\sigma_{2}^{i}(\tau(z))=\tau(\sigma_{1}(z))k_{i}$
for all $i\in\{0,\ldots,m-1\}$ and all $z\in K$. Since the $\sigma_{i}$ and $\tau$ commute, we know $\sigma_{2}^{i}(\tau(z))=\tau(\sigma_{2}^{i}(z))$. This implies
$k_{i}(\tau\big{(}\sigma_{2}^{i}(z)-\sigma_{1}(z)\big{)})=0$
for all $i\in\{0,\ldots,m-1\}$ and all $z\in K$, i.e.
(3.4) $k_{i}=0\text{ or }\sigma_{2}^{i}=\sigma_{1}$
for $i\in\{0,\ldots,m-1\}$. Since $\sigma_{1}=\sigma_{2}^{j}$ and $\sigma_{i}$ has order $n\geq m$ it follows that $H(t)=kt^{j}$ for some $k\in K$. Since $H$ is an isomorphism, we know $k\in K^{\times}$. ∎
We now prove our First Main Theorem:
Theorem 3.3.
Let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of $\mathrm{Gal}(K/F)$. Let $t^{m}-a_{1}\in K[t;\sigma_{1}]$ and $t^{m}-a_{2}\in K[t;\sigma_{2}]$ be not right-invariant, and suppose that $n\geq m$. Then
$K[t;\sigma_{1}]/K[t;\sigma_{1}](t^{m}-a_{1})\not\cong K[t;\sigma_{2}]/K[t;\sigma_{2}](t^{m}-a_{2})$
for any choice of $a_{i}\in K\smallsetminus F$.
Proof.
Since $\sigma_{1},\sigma_{2}$ both generate $\mathrm{Gal}(K/F)$, there is some $1<j<n$, with $(j,n)=1$, such that $\sigma_{1}=\sigma_{2}^{j}$. Let $\ell$ be the least positive integer such that $\ell j>m$; since $\sigma_{1}\neq\sigma_{2}$, $\ell<m$. Suppose that $H:K[t;\sigma_{1}]/K[t;\sigma_{1}](t^{m}-a_{1})\rightarrow K[t;\sigma_{2}]/K[t;\sigma_{2}](t^{m}-a_{2})$ is an $F$-algebra isomorphism. Then $H|_{K}=\tau\in{\rm Gal}(K/F)$ and $H(t)=kt^{j}$ for some $k\in K^{\times}$ by Lemma 3.2. Since the Galois group is cyclic, and thus abelian, $\tau$ commutes with $\sigma_{i}$ on $K$, for $i\in\{1,2\}$. Let us rename the indeterminants in both algebras to $t_{1}$ and $t_{2}$, respectively, for a clearer argument. For any $s\geq 1$, put $\gamma_{s}(k)=k\sigma_{2}^{j}(k)\sigma_{2}^{2j}(k)\cdots\sigma_{2}^{(s-1)j}(k)$. Then one can show by induction that
$H(t_{1}^{s})=\begin{cases}\gamma_{s}(k)t_{2}^{sj}&\text{if $1\leq s<\ell$;}\\ \gamma_{\ell}(k)\sigma^{\ell j-m}(a_{2})t_{2}^{\ell j-m}&\text{if $s=\ell$.}\end{cases}$
Since $1\leq\ell<m$, we have
$t_{1}\cdot t_{1}^{\ell}=t_{1}^{\ell}\cdot t_{1}=\begin{cases}t_{1}^{\ell+1}&\text{if $\ell+1<m$};\\ a_{1}&\text{if $\ell+1=m$.}\end{cases}$
However, we have
$H(t_{1}^{\ell})H(t_{1})=\gamma_{\ell}(k)\sigma^{\ell j-m}(a_{2})t_{2}^{\ell j-m}\cdot kt_{2}^{j}$
so that
$H(t_{1}^{\ell})H(t_{1})=\begin{cases}\gamma_{\ell+1}(k)\sigma^{\ell j-m}(a_{2})t_{2}^{\ell j+j-m}&\text{if $(\ell+1)j<2m$;}\\ \gamma_{\ell+1}(k)\sigma^{\ell j-m}(a_{2})\sigma^{{\ell j+j-2m}}(a_{2})t_{2}^{\ell j+j-2m}&\text{if $2m\leq(\ell+1)j$}\end{cases}$
whereas
$H(t_{1})H(t_{1}^{\ell})=kt_{2}^{j}\cdot\gamma_{\ell}(k)\sigma_{2}^{\ell j-m}(a_{2})t_{2}^{\ell j-m},$
so that
$H(t_{1})H(t_{1}^{\ell})=\begin{cases}\gamma_{\ell+1}(k)\sigma_{2}^{\ell j-m+j}(a_{2})t_{2}^{\ell j+j-m}&\text{if $(\ell+1)j<2m$;}\\ \gamma_{\ell+1}(k)\sigma_{2}^{\ell j-m+j}(a_{2})\sigma_{2}^{\ell j-2m+j}(a_{2})t_{2}^{\ell j+j-2m}&\text{if $2m\leq(\ell+1)j$}.\end{cases}$
Hence $H$ being well-defined implies that
$\sigma_{2}^{\ell j-m}(a_{2})=\sigma_{2}^{\ell j-m+j}(a_{2})$
i.e. that $a_{2}=\sigma_{2}^{j}(a_{2})=\sigma_{1}(a_{2})$. It follows that $H$ is well-defined only if $\sigma_{1}(a_{2})=a_{2}$, i.e. if $a_{2}\in{\rm Fix}(\sigma_{2})=F$. But this is impossible: $\sigma_{1}$ generates the cyclic Galois group, and $a_{2}\notin F$ by assumption. ∎
Corollary 3.4.
Let $n\geq m$. Let $\varphi$ be the Euler function. Then there are $\varphi(n)$ different classes of proper Petit algebras of type $K[t;\sigma_{i}]/K[t;\sigma_{i}](t^{m}-a)$; one class for each choice of generator $\sigma_{i}$ of $\mathrm{Gal}(K/F)$. Algebras in two different classes are not isomorphic.
The assumption that $n\geq m$ in Lemma 3.2 is indeed needed in the proof for (3.3). If $n<m$ we obtain a weaker version:
Lemma 3.5.
Let $F$ be an arbitrary field and let $n\geq 3$ be the degree of a cyclic Galois extension $K/F$. Let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of the Galois group $\mathrm{Gal}(K/F)$. Since $\sigma_{1},\sigma_{2}$ both generate $\mathrm{Gal}(K/F)$, there is some $1<j<n$, with $(j,n)=1$, such that $\sigma_{1}=\sigma_{2}^{j}$. Let $f\in K[t;\sigma_{1}]$ and $g\in K[t;\sigma_{2}]$ be both not right-invariant of degree $m$, $n<m$. Let
$H:K[t;\sigma_{1}]/K[t;\sigma_{1}]f\rightarrow K[t;\sigma_{2}]/K[t;\sigma_{2}]g$
be an $F$-algebra isomorphism between the two proper nonassociative algebras $K[t;\sigma_{1}]/K[t;\sigma_{1}]f$ and $K[t;\sigma_{2}]/K[t;\sigma_{2}]g$. Then
$H(t)=k_{j}t^{j}+k_{n+j}t^{n+j}+\ldots+k_{sn+j}t^{sn+j}$
for the largest integer $s$, such that $sn+j<m-1$ and $0\leq l\leq s$, where $k_{ln+j}\in K$. Thus
$H(\sum_{i=0}^{m-1}x_{i}t^{i})=(\sum_{i=0}^{m-1}\tau(x_{i})h(t)^{i}){\rm mod}_{r}g.$
Moreover, for
$h(t)=k_{j}t^{j}+k_{n+j}t^{n+1+j}+\ldots+k_{sn+j}t^{sn+j}$
we have $h(t)^{i}$ is well defined for all $i\leq m-1$, i.e., all $i$th powers of $h(t)$ are power-associative for $i\leq m-1$, and $h(t)h(t)^{m-1}=\sum_{i=0}^{m-1}\tau(a_{i})h(t)^{i}{\rm mod}_{r}g.$
Proof.
Again, $H$ restricts to some $\tau\in{\rm Gal}(K/F)$ and if we let $H(t)=\sum_{i=0}^{m-1}k_{i}t^{i}$ for some $k_{i}\in K$ then we obtain (3.4) for all $i\in\{0,\ldots,m-1\}$ and all $z\in K$, i.e. $k_{i}=0$ or $\sigma_{2}^{i}=\sigma_{2}^{j}$ for $i\in\{0,\ldots,m-1\}$, that is $k_{i}=0$ or $\sigma_{2}^{i-j}=\sigma_{2}$ for $i\in\{0,\ldots,m-1\}$. Now $\sigma_{2}^{i-j}=\sigma_{2}$ if and only if $i-j=nl$ for some $l\in\mathbb{Z}$, since $\sigma_{2}$ has order $n<m-1$. Therefore (3.4) implies $k_{i}=0$ for every $i\neq nl+j$, $l\in\mathbb{N}_{0}$, $i\in\{0,\ldots,m-1\}$. Thus
$H(t)=k_{j}t^{j}+k_{n+j}t^{n+1+j}+\ldots+k_{sn+j}t^{sn+j}$
for some integer $s$, $sn+j<m-1$. Put $h(t)=k_{j}t^{j}+k_{n+j}t^{n+j}+\ldots+k_{sn+j}t^{sn+j}$.
Now $H(t^{i})=H(t)^{i}=g(t)^{i}{\rm mod}_{r}g$ is well defined for all $i\in\{0,\ldots,m-1\}$, since so is $t^{i}$, and
$H(t)H(t^{m-1})=h(t)h(t)^{m-1}{\rm mod}_{r}g=H(t^{m})=\sum_{i=0}^{m-1}\tau(a_{i})h(t)^{i}{\rm mod}_{r}g.$
4. Counting the non-isotopic nonassociative cyclic algebras of prime degree
Two semifields $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma,a)$ and $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma,b)$ are isomorphic if and only if they are isotopic [San1962, Theorem 9]. The proof of [San1962, Theorem 9] can be generalized and yields our Second Main Theorem:
Theorem 4.1.
Let $K/F$ be a cyclic Galois field extension, $m\geq 3$, and $\sigma_{1}$, $\sigma_{2}$ two generators of ${\rm Gal}(K/F)$. Let $(K/F,\sigma_{1},a_{1})$ and $(K/F,\sigma_{2},a_{2})$ be two proper nonassociative algebras. If $(K/F,\sigma_{1},a_{1})$ and $(K/F,\sigma_{2},a_{2})$ are isotopic, then $\sigma_{1}=\sigma_{2}$.
We break up the proof of this result and start with a generalization of [San1962, Theorem 4] which does not require that the algebras are division algebras:
Theorem 4.2.
(i) Let $(Q,P,U)$ be an isotopy between two proper nonassociative algebras $A_{1}$ and $A_{2}$ over $F$ with middle nucleus $K$. Then for every $a\in{\rm Nuc}_{m}(A_{2})=K$ there is $c\in{\rm Nuc}_{m}(A_{1})=K$ such that
$PL_{a}=L_{c}P.$
(ii) Let $K/F$ be a Galois field extension of degree $n\geq 3$, and $\sigma_{1}$, $\sigma_{2}$ two generators of ${\rm Gal}(K/F)$. Suppose that $m\leq n$. Let $(Q,P,U)$ be an isotopy between two proper nonassociative algebras $K[t;\sigma_{1}]/K[t;\sigma_{1}](t^{m}-a_{1})$ and $K[t;\sigma_{2}]/K[t;\sigma_{2}](t^{m}-a_{2})$. Then there exists some $V\in{\rm Aut}(K)$ and some $h_{i}\in K^{\times}$ such that
$P=\left[\begin{array}[]{cccccc}VR(h_{1})&&...&\\ &VR(h_{2})&...&\\ &&VR(h_{3})&\\ ...&...&...&...\\ &&...&\\ &&...&VR(h_{m})\\ \end{array}\right].$
Proof.
As in [San1962] we can show that for any $a\in{\rm Nuc}_{m}(A_{2})=K$ there is $c\in{\rm Nuc}_{m}(A_{1})=K$ such that $PL_{a}=L_{c}P$ where $L_{c}$ is the left multiplication in $A_{1}$ and $L_{a}$ the one in $A_{2}$.
(ii) We continue as in [San1962, p. 244], using (i) to compute $P$ to be
$P=\left[\begin{array}[]{cccccc}VR(h_{11})&\sigma_{2}VR(h_{12})&...&\sigma_{2}^{m-1}VR(h_{1m})\\ \sigma_{2}^{-1}VR(h_{21})&VR(h_{22})&...&\sigma_{2}^{m-2}VR(h_{2m})\\ \sigma_{2}^{-2}VR(h_{31})&&VR(h_{33})&\\ ...&...&...&...\\ &&...&\\ \sigma_{2}^{-m+1}VR(h_{m1})&\sigma_{2}^{-m+2}VR(h_{m2})&...&VR(h_{mm})\\ \end{array}\right]$
for some $V\in{\rm Aut}(K)$ where $R(h_{ij})$ is right multiplication with $h_{ij}\in K$, as in [San1962, (34)]. Let $Q(1)=u_{0}+u_{1}t+\dots+t^{m-1}u_{m-1}$ for some $u_{i}\in K$ and $Q(t)=a_{0}+a_{1}t+\dots+t^{m-1}a_{m-1}$ for some $a_{i}\in K$. Analogously as in [San1962, (19)], we have $PL_{Q(z)}=L_{z}PL_{Q(1)}$. Put $z=t$ in $PL_{Q(z)}=L_{z}PL_{Q(1)}$, this way e obtain a set of linear equations identical to the ones in [San1962, (31)], just with the formula in the left hand side employing $\sigma_{1}$ and the ones on the right hand side $\sigma_{2}$. Arguing as in [San1962, p. 246], if exactly one element in the first row of $P$ is non-zero, then exactly one element in each row or column of $P$ is non-zero, and these will all lie on a transversal. We obtain hence analogously as in [San1962, p. 248] that
$P=\left[\begin{array}[]{cccccc}VR(h_{11})&&...&\\ &VR(h_{22})&...&\\ &&VR(h_{33})&\\ ...&...&...&...\\ &&...&\\ &&...&VR(h_{mm})\\ \end{array}\right]$
for some $h_{ij}\in K^{\times}$, call them $h_{i}=h_{ii}$.
We can now do the proof of Theorem 4.1, which is the case that $m=n$:
Proof.
Suppose that the algebras are isotopic. Let $(Q,P,U)$ be such an isotopy. By Theorem 4.2,
$P=\left[\begin{array}[]{cccccc}VR(h_{1})&&...&\\ &VR(h_{2})&...&\\ &&VR(h_{3})&\\ ...&...&...&...\\ &&...&\\ &&...&VR(h_{m})\\ \end{array}\right].$
Without loss of generality we assume that $h_{1}=1$. This is possible, since for every $a\in K^{\times}$, $(id,D_{a},D_{a})$ is an autotopism of $A_{i}$, where $D_{a}$ denotes the diagonal matrix with $a$ as its diagonal entries. Hence we can multiply $P$ by a suitable autotopism to achieve $h_{1}=1$. As in [San1962, (44)], we also obtain that $Q=U=PL_{Q(1)}$. We continue along the lines of Sandler’s proof to see that also in our setting, $h_{k}=h_{1}\sigma_{2}(h_{1})\sigma_{2}^{2}(h_{1})\cdots\sigma_{2}^{k-1}(h_{1})$. Note that up to here, it is not required that the algebra is a division algebra. Then a short calculation analogous to the one given in [San1962, p. 264] shows that we obtain either that $\sigma(V(a_{1}))=a_{2}h_{1}\sigma_{2}(h)\sigma_{2}^{2}(h)\cdots\sigma_{2}^{m-1}(h)=a_{2}N_{K/F}(h)$ or $V(a_{1})=a_{2}h_{1}\sigma_{2}(h)\sigma_{2}^{2}(h)\cdots\sigma_{2}^{m-1}(h)=a_{2}N_{K/F}(h)$ for some $h\in K^{\times}$. In either case, this implies that $P$ is an isomorphism between the two algebras by Theorem 2.3. This means that $\sigma_{1}=\sigma_{2}$ by Theorem LABEL:sigmadistinct. ∎
Note that the above proof does not require that the field $F$ is finite.
For a finite field $F=\mathbb{F}_{q}$ and $t^{m}-a_{1}\in\mathbb{F}_{q^{n}}[t;\sigma_{1}]$, $t^{m}-a_{2}\in\mathbb{F}_{q^{n}}[t;\sigma_{2}]$ both irreducible and not right-invariant, these algebras are the dual algebras of Sandler’s semifields. For $n=m$ we obtain from [San1962, Theorem 9], Theorem 4.1:
Corollary 4.3.
Let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of $\mathrm{Gal}(\mathbb{F}_{q^{m}}/\mathbb{F}_{q})$. Let $t^{m}-a_{1}\in\mathbb{F}_{q^{m}}[t;\sigma_{1}]$ and $t^{m}-a_{2}\in\mathbb{F}_{q^{m}}[t;\sigma_{2}]$ be irreducible and not right-invariant. Then
$\mathbb{F}_{q^{m}}[t;\sigma_{1}]/\mathbb{F}_{q^{m}}[t;\sigma_{1}](t^{m}-a_{1})$
and
$\mathbb{F}_{q^{m}}[t;\sigma_{2}]/\mathbb{F}_{q^{m}}[t;\sigma_{2}](t^{m}-a_{2})$
are non-isomorphic (respectively, non-isotopic) semifields.
Theorem 4.4.
([BrPumSt2018, Theorem 3.1]) Let $m$ be prime. Fix a generator $\sigma$ of ${\rm Gal}(\mathbb{F}_{q^{m}}/\mathbb{F}_{q})$.
(i) If $m$ does not divide $q-1$ then there are exactly
$\frac{q^{m}-q}{m(q-1)}$
non-isomorphic semifields $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma,a)$ of order $q^{m^{2}}$.
(ii) If $m$ divides $q-1$ then there are exactly
$m-1+\frac{q^{m}-q-(q-1)(m-1)}{m(q-1)}$
non-isomorphic semifields $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma,a)$ of order $q^{m^{2}}$.
Corollary 4.5.
(i) Let $m$ not be prime and let $\varphi(m)$ be the number of generators of the cyclic Galois group ${\rm Gal}(K/F)$ of order $n$. Then we have $\varphi(m)$ different classes of non-isotopic semifields $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{i},a)$, one class for each choice of generator $\sigma_{i}$ of ${\rm Gal}(K/F)$.
(ii) If $m$ is prime and does not divide $q-1$, then there are exactly
$\frac{(q^{m}-q)(m-1)}{(q-1)m}$
non-isotopic Sandler semifields of order $q^{m^{2}}$ with center $F=\mathbb{F}_{q}$ and left, middle and right nucleus $K=\mathbb{F}_{q^{m}}$.
(iii) If $m$ is prime and divides $q-1$, then there are exactly
$(m-1)^{2}+(m-1)\frac{q^{m}-q-(q-1)(m-1)}{m(q-1)}$
non-isotopic Sandler semifields of order $q^{m^{2}}$ with center $F=\mathbb{F}_{q}$ and left, middle and right nucleus $K=\mathbb{F}_{q^{m}}$.
We deduce that there are more non-isomorphic Sandler semifields than previously counted in [BrPumSt2018, Theorem 3.1]:
Theorem 4.6.
(i) Let $\varphi(m)$ be the number of generators of the cyclic Galois group ${\rm Gal}(\mathbb{F}_{q^{m}}/\mathbb{F}_{q})$ of order $m$ ($\varphi$ being the Euler function). Then we have $\varphi(m)$ different classes of non-isomorphic semifields $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{i},a)$, one class for each choice of generator $\sigma_{i}$ of ${\rm Gal}(\mathbb{F}_{q^{m}}/\mathbb{F}_{q})$.
(ii) If $m$ is prime and $m$ does not divide $q-1$, then there are exactly
$\frac{(m-1)(q^{m}-q)}{m(q-1)}$
non-isomorphic Sandler semifields of order $q^{m^{2}}$ with center $\mathbb{F}_{q}$ and left, middle and right nucleus $\mathbb{F}_{q^{m}}$.
(iii) If $m$ is prime and divides $q-1$, then there are exactly
$(m-1)((m-1)+\frac{q^{m}-q-(q-1)(m-1)}{m(q-1)})$
non-isomorphic Sandler semifields of order $q^{m^{2}}$ with center $F=\mathbb{F}_{q}$ and left, middle and right nucleus $K=\mathbb{F}_{q^{m}}$.
Proof.
(i) is trivial.
(ii) For each $i$, $11\leq i\leq m-$ 1, there are exactly
$\frac{q^{m}-q}{m(q-1)}$
non-isomorphic Sandler semifields of order $q^{m^{2}}$ that are the opposite algebras of $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma^{i},a)$.
(iii) For each $i$, $11\leq i\leq m-$ 1 there are exactly
$m-1+\frac{q^{m}-q-(q-1)(m-1)}{m(q-1)}$
non-isomorphic Sandler semifields of order $q^{m^{2}}$ that are the opposite algebras of $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma^{i},a)$. In total there are therefore up to isomorphism exactly
$(m-1)(m-1+\frac{q^{m}-q-(q-1)(m-1)}{m(q-1)})=(m-1)^{2}+(m-1)\frac{q^{m}-q-(q-1)(m-1)}{m(q-1)}$
Sandler semifields of order $q^{m^{2}}$ with center $\mathbb{F}_{q}$ and left, middle and right nucleus $\mathbb{F}_{q^{m}}$.
When $m$ is not prime, we have the following statement, again for $m=n$.
Theorem 4.7.
[Ste2014, Theorem 4.4, Corollary 4.5] Suppose that $(K/F,\sigma,a)$ is a nonassociative cyclic algebra where $K/F$ is of degree $m$. If $a$ does not lie in a proper subfield of $K/F$, then $(K/F,\sigma,a)$ is a division algebra.
Example 4.8.
We briefly discuss the $n=4$: Let $F$ be a finite field of odd prime characteristic. Let $K$ be the cyclic Galois field extension of degree four of $F$ with ${\rm Gal}(K/F)=\langle\sigma\rangle$. Then $K$ has exactly one quadratic subfield $E={\rm Fix}(\sigma^{2})$.
Let $a\in K\smallsetminus F$ and consider $(K/F,\sigma,a)$. This is a 16-dimensional algebra over $F$. By Theorem 4.7, if $a\notin E$ then $(K/F,\sigma,a)$ is a division algebra.
Note that the nonassociative cyclic $E$-algebra $B=(K/E,\sigma^{2},a)$ is an $F$-subalgebra of $(K/F,\sigma,a)$. If $a\not\in E$, then $B$ is a proper nonassociative quaternion division algebra over $E$.
If $a\in E$, then $B$ is the right nucleus of $A$. In this case, $B$ is an associative quaternion algebra over $E$, and this is a never a division algebra over a finite field. Since $B$ is a division algebra if and only if $(K/F,\sigma,a)$ is a division algebra, when $F$ is a finite field this case does not yield a semifield.
We thus showed by a direct proof that $(K/F,\sigma,a)$ is a semifield if and only if $a\not\in E$. So the first step is to count all these $a$’s.
5. An explicit parametrization of nonassociative cyclic algebras in odd characteristic
Two proper semifields coordinatize the same non-Desarguesian plane if and only if they are isotopic. The subsequent results thus yield a full parametrization of the non-Desarguesian planes related to the semifields $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma,a)$.
5.1. When $m$ is an odd prime
Suppose that $m$ is an odd prime and that $F=\mathbb{F}_{q}$ has odd characteristic and contains a primitive $m$th root of unity $\zeta\in F$, which implies that $gcd(m,p)=1$. Let $K=\mathbb{F}_{q^{m}}$ and fix a choice of generator $\sigma$ of $\mathrm{Gal}(\mathbb{F}_{q^{m}}/\mathbb{F}_{q})$. Since $\mathbb{F}_{q^{m}}$ is the splitting field of some polynomial $x^{m}-b$, with $b\notin(\mathbb{F}_{q}^{\times})^{m}$, we may choose $\beta\in\mathbb{F}_{q^{m}}$ to be a root of $x^{m}-b$ that satisfies $\sigma(\beta)=\zeta\beta$. Then
$\{1,\beta,\beta^{2},\ldots,\beta^{m-1}\}$
is an $\mathbb{F}_{q}$-basis for $\mathbb{F}_{q^{m}}$ and $\sigma(\beta^{k})=\zeta^{k}\beta^{k}$ for all $0\leq k<m$.
We now parametrize the nonassociative cyclic algebras $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma,a)$ of odd prime degree $m$, applying results from [NevPu2024].
Let $\mathcal{I}=\mathcal{P}(\{0,1,\ldots,m-1\})\smallsetminus\{\{0\},\emptyset\}$ be the set of nonempty subsets of $\{0,1,\ldots,m-1\}$, excluding the set $\{0\}$. For each $I\in\mathcal{I}$, define
$K(I)=\{\sum_{i=0}^{m-1}a_{i}\beta^{i}\mid\text{ for all }i\in I,a_{i}\in\mathbb{F}_{q}^{\times},\text{and for all }i\notin I,a_{i}=0\}\subset\mathbb{F}_{q^{m}}\smallsetminus\mathbb{F}_{q}.$
Let $I\in\mathcal{I}$ be such that $|I|=k+1\geq 2$. Let $\mathbb{F}_{q}^{[\times k]}=\mathbb{F}_{q}^{\times}\times\mathbb{F}_{q}^{\times}\times\cdots\times\mathbb{F}_{q}^{\times}$ be the $k$-fold direct product. If the elements of $I$ are $i_{0}<i_{1}<\cdots<i_{k}$ then let
[…]
Write $\mathbb{F}_{q}^{[\times k]}/\Delta_{I}$ for any fixed choice of representatives for these cosets. For any fixed choice of $a_{i_{0}}\in F^{\times}$, define
$K(I;a_{i_{0}})=\{a_{i_{0}}\beta^{i_{0}}+\sum_{j=1}^{k}a_{i_{j}}\beta^{i_{j}}:(a_{i_{1}},a_{i_{2}},\ldots,a_{i_{k}})\in\mathbb{F}_{q}^{[\times k]}/\Delta_{I}\}.$
Finally, set $K(I;a_{i_{0}})=\{a_{i_{0}}\beta^{i_{0}}\}$ when $I=\{i_{0}\}\in\mathcal{I}$ and define
$\mathcal{S}(\mathbb{F}_{q^{m}})=\bigcup_{I\in\mathcal{I}}K(I;1)\subset\mathbb{F}_{q^{m}}\smallsetminus\mathbb{F}_{q}.$
Theorem 5.1.
Suppose $m$ is an odd prime and $\mathbb{F}_{q}$ contains a primitive $m$th root of unity $\zeta$. Then the distinct isomorphism classes of nonassociative cyclic algebras of degree $m$ over $F$ are represented by
$(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma,a)$
where:
  • $\mathbb{F}_{q^{m}}$ is the unique cyclic Galois field extensions of $\mathbb{F}_{q}$ of degree $m$;
  • $\sigma$ is a generator of $\mathrm{Gal}(\mathbb{F}_{q^{m}}/\mathbb{F}_{q})$;
  • $a\in\mathcal{S}(\mathbb{F}_{q^{m}})$.
This is [NevPu2024, Theorem 5.2] applied to $F=\mathbb{F}_{q}$. Note that for finite fields, the norm map is surjective, hence $N_{K/F}(K^{\times})=F^{\times}$ and so $F^{\times}/N_{K/F}(K^{\times})=\{1\}$ is trivial.
5.2. When $m=2$
Suppose that $m=2$ and that $\mathbb{F}_{q}$ has odd characteristic. We obtain from [NevPu2024, Theorem 4.2, Corollary 4.3]:
Theorem 5.2.
Suppose $\mathbb{F}_{q^{2}}=\mathbb{F}_{q}(\sqrt{c})$ for some $c\in\mathbb{F}_{q}^{\times}\smallsetminus(\mathbb{F}_{q}^{\times})^{2}$ and let $\sigma\in\mathrm{Gal}(\mathbb{F}_{q^{2}}/\mathbb{F}_{q})$ be nontrivial. Then the distinct isomorphism classes of the nonassociative quaternion algebras with nucleus $\mathbb{F}_{q^{2}}$ are represented by $(\mathbb{F}_{q^{2}}/\mathbb{F}_{q},\sigma,a)$, where $a$ is chosen from the set
$\mathcal{S}_{2}(\mathbb{F}_{q^{2}})=\{\sqrt{c},1+s\sqrt{c}\mid s\in\mathbb{F}_{q}^{\times}/\{\pm 1\}\}.$
Alternatively, we can parametrize these distinct isomorphism classes by elements of the set
$\mathcal{S}^{\prime}(\mathbb{F}_{q^{2}})=\{t+\sqrt{c}\mid t\in\{0\}\cup\mathbb{F}_{q}^{\times}/\{\pm 1\}\}.$
6. What does this mean for the related linear codes?
Our results raise several questions that will be addressed in a future paper. It would certainly be interesting to see in how far our Second Main Theorem can be generalized to hold for other Petit algebras $K[t;\sigma]/K[t;\sigma]f$, when we still assume that the degree of $f$ is less or equal to the order $n$ of the automorphism $\sigma$ (the order may be infinite), but where $K/F$ does not need to be a Galois field extension. A straightforward generalization of the proof of Theorem 3.3 already shows:
Theorem 6.1.
Let $F$ be an arbitrary field and let $n\geq 3$ be the degree of the cyclic Galois extension $K/F$. Let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of $\mathrm{Gal}(K/F)$, and suppose that $n\geq m$. Let
$f(t)=t^{m}-\sum_{i=0}^{m-1}a_{i}t^{i}\in K[t;\sigma_{1}],\quad g(t)=t^{m}-b\in K[t;\sigma_{2}]$
such that all $a_{i}\in K\smallsetminus F$, and assume that $f$ and $g$ both are not right-invariant (i.e. the two Petit algebras are proper nonassociative algebras. Then
$K[t;\sigma_{1}]/K[t;\sigma_{1}]f\not\cong K[t;\sigma_{2}]/K[t;\sigma_{2}](t^{m}-b).$
We expect to obtain similar results for large classes of Petit algebras $K[t;\sigma_{1}]/K[t;\sigma_{1}]f$.
6.1. $(f,\sigma,\delta)$-codes
Skew polynomials $f\in R=K[t;\sigma,\delta]$ can be used to build linear codes (for a comprehensive exposition, cf. [BouLe2013]). Module $\sigma$-codes and $(f,\sigma,\delta)$-codes can be associated to the left ideals of a nonassociative Petit algebra $R/Rf$ for a suitable reducible $f\in R$ (choosing $\delta=0$ for the module $\sigma$-codes): the cyclic submodules studied in [BouUl2009, BUl2014, BUl2009] are exactly the left ideals in the Petit algebra $R/Rf$, the $(\sigma,\delta)$-codes of [BouLe2013] are the codes $\mathcal{C}$ associated to some left ideal of the Petit $R/Rf$. These ideals are generated by a right divisor $g$ of $f$.
If $\mathcal{C}$ is a linear code over $\mathbb{F}_{q}$ of length $m$, then $\mathcal{C}$ is a $\sigma$-constacyclic code with constant $d$ if and only if the skew-polynomial representation $\mathcal{C}(t)$ with elements $a(t)$ obtained from $(a_{0},\dots,a_{m-1})\in\mathcal{C}$ is a left ideal of the Petit algebra $\mathbb{F}_{q}[t;\sigma]/\mathbb{F}_{q}[t;\sigma](t^{m}-d)$, which is generated by a monic right divisor $g$ of $f$ in $R$.
Let $f(t)=t^{m}-d\in R=K[t;\sigma]$ and $F={\rm Fix}(\sigma)$. Then any right divisor $g$ of degree $k$ of $f=t^{m}-d$ can be used to construct a $\sigma$-constacyclic $[m,m-k]$-code with constant $d$.
Let $(K/F,\sigma,d)$ be a nonassociative cyclic division algebra of prime degree $m$, i.e. $f$ is irreducible. Therefore there are no $\sigma$-constacyclic codes with constant $d$, apart from the $[m,m]$-code associated with the nonassociative cyclic algebra $(K/F,\sigma,d)$ itself. So we need to look at algebras that are not division algebras in this context.
Let $K=\mathbb{F}_{q^{n}}[t;\sigma]$ and $a_{i}\in K\smallsetminus F$. If $n\geq 3$ and $n>m$ then the Petit algebras $\mathbb{F}_{q}[t;\sigma]/\mathbb{F}_{q}[t;\sigma](t^{m}-a_{i})$ are nonassociative. Since
$K[t;\sigma_{1}]/K[t;\sigma_{1}](t^{m}-b)\not\cong K[t;\sigma_{2}]/K[t;\sigma_{2}](t^{m}-b)$
for any two distinct generators $\sigma_{1}$ and $\sigma_{2}$ of the Galois group $\mathrm{Gal}(K/F)$ and any choice of $a_{i}\in K\smallsetminus F$, we expect that the choice of generator $\sigma_{i}$ will also affect the $\sigma_{i}$-constacyclic codes we obtain, since these correspond to the left ideals of two non-isomorphic algebras.
6.2. MRD-codes
Linear maximum rank distance codes (MRD-codes) can be obtained from skew polynomial rings $K[t;\sigma]$, employing the matrices representing the left multiplication in a division algebra $K[t;\sigma]/K[t;\sigma]f$, with entries in $K$.
Isotopy classes of finite semifields of dimension $n$ over $F=\mathbb{F}_{q}$ are in one-one correspondence with equivalence classes of $\mathbb{F}_{q}$-linear MRD-codes in $M_{n}(\mathbb{F}_{q})$ with minimum distance $n$. Over finite fields, Sandler showed that two algebras $K[t,\sigma]/K[t,\sigma](t^{m}-a)$ and $K[t,\sigma]/K[t,\sigma](t^{m}-b)$ are isotopic iff they are isomorphic.
Our results show that for $n\geq m$, the MRD codes obtained from the Petit algebras $K[t,\sigma]/K[t,\sigma](t^{m}-a)$ depend on the choice of the generator of the Galois group [She2019]: Different choices of generators of ${\rm Gal}(K/F)$ yield non-equivalent $F$-linear MRD-codes.
Since these codes are special cases of the codes constructed in [She2019] we conjecture that the equivalence classes of all of Sheekey’s codes also will depend on the choice of generator of the Galois group involved.
Moreover, we achieved an explicit parametrization of all semifields that are nonassociative cyclic algebras of a fixed prime degree over a fixed base field $F=\mathbb{F}_{q}$ which contains a primitive $n$th root of unity. That means we computed the number of non-isotopic semifields of order $q^{m^{2}}$ obtained via the construction $(K/F,\sigma,a)$ when ${\rm Fix}(\sigma)=F=\mathbb{F}_{q}$, and gave one representative for each isotopy class. Isotopy classes of finite semifields of dimension $m$ over $F=\mathbb{F}_{q}$ are in one-one correspondence with equivalence classes of $\mathbb{F}_{q}$-linear MRD-codes in $M_{m}(\mathbb{F}_{q})$ with minimum distance $m$. Thus we now have one representative for each corresponding MRD-code.